# AutoFT: Robust Fine-Tuning by Optimizing Hyperparameters on OOD Data

### 18 Jan 2024 v1

@@ 6 Results
Table 5: In addition to enhancing robustness, AutoFT can improve performance in IID settings, where there is no distribution shift. Here, AutoFT uses ID validation examples to optimize hyperparameters. AutoFT achieves the highest ID performance on CalTech-101 and matches the best baseline on Stanford Cars.
MNIST CIFAR10.2 CIFAR10.1
Method ID OOD ID OOD ID OOD
FT 99.8 62.3 97.9 96.2 98.3 93.4
AutoFT 99.8 61.1 98.2 96.3 98.4 93.4
Table 6: Performance comparison of models fine-tuned on new distributions with hyperparameters learned by AutoFT on a different dataset, CIFAR-10. Hyperparameters transfer better to datasets that are more similar to the original fine-tuning distribution.
Validation Dataset CIFAR-10.1 CIFAR-10.2
CIFAR-10-C 97.4 (0.2) 93.5 (0.2)
CIFAR-10 96.9 (0.2) 93.2 (0.1)
CIFAR-10.1 96.8 (0.3) 92.8 (0.2)
CIFAR-10.2 96.7 (0.5) 93.8 (0.0)
Table 7: Performance of AutoFT with different validation datasets for hyperparameter evaluation. All AutoFT runs above optimize hyperparameters for fine-tuning on CIFAR-10. AutoFT produces a robust fine-tuned model with all validation datasets, with the best results achieved when the validation set is from a different distribution than the fine-tuning dataset, CIFAR-10.
6.1 Evaluation Under Distribution Shifts
Improvements on ImageNet, WILDS, and CIFAR distribution shifts. We evaluate AutoFT on nine natural distribution shifts: five from ImageNet, WILDS-FMoW, WILDS-iWildCam, and two subtle CIFAR-10 distribution shifts. For the OOD validation sets, we use up to $1000$ examples from ImageNet-C, the OOD validation splits of WILDS-FMoW and WILDS-iWildCam, and CIFAR-10-C, respectively. Tab. 3 contains results on ImageNet, WILDS-FMoW, and WILDS-iWildCam; results on CIFAR-10 are in  Tab. 9. AutoFT consistently outperforms all baselines on novel OOD distributions. These gains in OOD performance are maintained when ensembling zero-shot and fine-tuned models, following Wortsman et al. [92]. We report weight ensembling results with the mixture coefficient that yields the highest ID validation accuracy.
State-of-the-art performance on WILDS-iWildCam and WILDS-FMoW. To assess whether performance gains by AutoFT continue to hold on larger foundation models, we evaluate AutoFT with a ViT-L/14@336px model. As shown in Tabs. 1 and 2, AutoFT achieves significant gains of $6.0\%$ and $3.6\%$ in OOD and ID macro-F1, respectively, over the current leader on the WILDS-iWildCam benchmark [47], FLYP [30]. On the WILDS-FMoW benchmark, AutoFT also outperforms the current leader, SGD (Freeze-Embed) [51], by $1.5\%$ and $5.3\%$ in OOD and ID worst-region accuracy, respectively. On both benchmarks, AutoFT additionally outperforms the compute-intensive Model Soups [91], which ensembles more than 70 models fine-tuned with LP-FT and different augmentations and hyperparameters. AutoFT also outperforms LP-FT, a state-of-the-art baseline for fine-tuning.
6.2 Few-Shot Classification
In many real-world applications, limited amounts of labeled, task-specific data are available for fine-tuning. Few-shot classification thus serves as an important benchmark for evaluating the utility of fine-tuning approaches. Few-shot binary classification is a particularly challenging task for adaptation, given the small number of training examples. We evaluate on 4, 16, and 32 shot binary classification tasks from the PatchCamelyon and Rendered-SST2 datasets, following Radford et al. [72] and Goyal et al. [30]. PatchCamelyon contains digital pathology images for the detection of metastatic tissue. Rendered-SST2 focuses on optical character recognition for classifying text sentiment as positive or negative.
AutoFT demonstrates strong generalization capabilities with limited data, outperforming all baselines on all few-shot tasks in Tab. 4. For example, AutoFT outperforms FLYP by $3.7\%$ and full fine-tuning by $3.9\%$ in a challenging 4-shot classification task on Rendered-SST2.
6.3 Transfer Learning
In standard transfer learning settings, where no distribution shifts exist between train and test data, AutoFT shows strong performance. The primary goal here is to achieve strong in-distribution (ID) performance on the test set. In this setting, AutoFT uses a small ID validation set to optimize hyperparameters.
We compare AutoFT against other baselines on several transfer learning datasets: CalTech-101, StanfordCars, and Flowers-102 in Tab. 5. AutoFT matches or outperforms all baselines on 2 of the 3 transfer learning datasets, outperforming FLYP by $1.4\%$ on CalTech-101.
6.4 Ablation Studies
Hyperparameter transfer. Are the hyperparameters learned by AutoFT for one fine-tuning distribution effective when applied to new fine-tuning distributions? We first train AutoFT on CIFAR-10, using $100$ examples from corrupted CIFAR-10 for hyperparameter optimization. We then fine-tune a model with these learned hyperparameters on three new distributions: MNIST, CIFAR-10.1, and CIFAR-10.2. Finally, we evaluate the fine-tuned model’s performance on both the fine-tuning distribution and several distribution shifts. For MNIST, we evaluate the fine-tuned model’s performance on three distribution shifts: corrupted MNIST, EMNIST, and rotated MNIST. Similarly, we evaluate the model fine-tuned on CIFAR-10.1 on CIFAR-10.2, and vice versa. Further experimental details are in Appendix A. Results in Tab. 6 show model performance after fine-tuning with these transferred hyperparameters. On CIFAR-10.1 and CIFAR-10.2, AutoFT outperforms standard fine-tuning on ID and OOD accuracy. However, on MNIST, AutoFT results in slightly worse OOD performance than fine-tuning. This suggests that the hyperparameters optimized by AutoFT transfer more effectively to datasets that are similar to the original fine-tuning distribution.
Effect of the hyperparameter optimization dataset. How does the choice of dataset for hyperparameter optimization affect the performance of the fine-tuned model? AutoFT optimizes hyperparameters using a validation set distinct from both the ID fine-tuning distribution and OOD test sets. In Table 7, we assess the performance of AutoFT using different datasets for hyperparameter optimization.
The results suggest that AutoFT yields strong OOD performance across a range of validation datasets, with best OOD performance achieved when the validation set used for hyperparameter optimization is different from the ID dataset.
7 Conclusion
We introduce AutoFT, a novel data-driven approach for robust fine-tuning that optimizes an expressive set of hyperparameters using a small OOD validation set. AutoFT only requires a small amount of data from one naturally occurring OOD distribution—such data is often readily available or is possible to gather at a similar cost to that of the original ID training data. Our empirical results demonstrate that AutoFT consistently outperforms existing approaches and achieves a new state-of-the-art for robust fine-tuning. We hope that our work will inspire future research on data-driven approaches for robust fine-tuning.
Future directions. While our experiments show strong results in image classification, we have not yet verified the effectiveness of AutoFT outside this problem setting. Future research could explore extending the general workflow of AutoFT to other areas such as image segmentation, object detection, and natural language processing (NLP). Moreover, investigating alternative hyperparameter search spaces within AutoFT could offer further insights. Lastly, exploring the possibility of identifying a single set of hyperparameters that works effectively across multiple fine-tuning distributions presents an interesting direction for future work.
8 Acknowledgements
We thank Kyle Hsu, Lukas Haas, and other members of the IRIS lab for helpful feedback and discussions. We also thank Sachin Goyal for help with ImageNet experiments. This work was supported by KFAS and NSF. A.R. is supported by Schmidt Futures, Google, Apple, and Open Philanthropy. C.F. is supported by ONR grant N00014-21-1-2685 and the NSF CAREER award.
Appendix A Additional Experiments
A.1 ImageNet Distribution Shifts
A.1.1 Main Results
We provide detailed results for each ImageNet-derived distribution shift in Tab. 8. AutoFT outperforms all baselines on 9 out of 10 evaluation settings. FLYP, the prior state-of-the-art, has strong ID accuracy: $82.6\%$ and $82.9\%$ without and with ensembling, whereas AutoFT has $81.8\%$ and $82.4\%$. Despite its high ID accuracy, the model fine-tuned with FLYP has substantially worse OOD accuracy than AutoFT: $60.2\%$ and $63.2\%$ without and with ensembling, compared to $61.5\%$ and $64.3\%$ for AutoFT. We observe a similar effect with all other points of comparison, indicating that these methods for robust fine-tuning were “overfitting” to the ID distribution in a way that AutoFT does not.
Without Ensembling With Ensembling Methods ID Im-V2 Im-R Im-A Sketch ObjectNet Avg. OOD ID Im-V2 Im-R Im-A Sketch ObjectNet Avg. OOD Zeroshot 68.3 61.9 77.7 50.0 48.3 55.4 58.7 68.3 61.9 77.7 50.0 48.3 55.4 58.7 LP 79.9 69.8 70.8 46.4 46.9 52.1 57.2 80.0 70.3 72.4 47.8 48.1 52.8 58.3 FT 81.3 71.2 66.1 37.8 46.1 53.3 54.9 82.5 72.8 74.9 48.1 51.9 59.0 61.3 L2-SP 81.7 71.8 70.0 42.5 48.5 56.2 57.8 82.2 72.9 75.1 48.6 51.4 58.9 61.4 LP-FT 81.7 72.1 73.5 47.6 50.3 58.2 60.3 82.1 72.8 75.3 50.1 51.7 59.2 61.8 FLYP 82.6 73.0 71.4 48.1 49.6 58.7 60.2 82.9 73.5 76.0 53.0 52.3 60.8 63.1 AutoFT 81.8 73.1 72.4 48.8 49.8 63.5 61.5 82.4 73.6 76.4 53.1 52.6 65.6 64.3
Table 8: AutoFT consistently outperforms all baselines on each ImageNet distribution shift.
A.1.2 Few-Shot Classification
Figure 5 shows the average OOD accuracy on five ImageNet-derived distribution shifts against ID accuracy on ImageNet for 4, 16, and 32 shot classification. AutoFT outperforms baselines in effective robustness in all three settings. In 32-shot classification, AutoFT improves average OOD accuracy by $2.2\%$ compared to the leading baseline, FLYP.
Figure 5: AutoFT enhances effective robustness over existing methods, indicated by vertical distance. AutoFT results in better OOD accuracy at a given ID accuracy across varying degrees of weight interpolation between the fine-tuned and zero-shot models.
A.2 CIFAR Distribution Shifts
@@
Table 9: AutoFT outperforms fine-tuning by $2.2\%$ on CIFAR-10.2 and by $1.4\%$ on CIFAR-10.1, using only 100 samples from CIFAR-10-C. AutoFT additionally outperforms WiSE-FT with weight ensembling.
AutoFT outperforms prior approaches on the subtle CIFAR-10.1 and CIFAR-10.2 distribution shifts.
A.3 Expressivity of Hyperparameter Space
@@
Table 10: AutoFT yields better ID and OOD performance than its variant with per-layer learning rates, weight decays, and L1/L2 distances to pretrained parameters, indicating that more expressive hyperparameter spaces overfit to the validation set, degrading performance.
@@
We investigate the effect of expressivity of the hyperparameter search space. We compare AutoFT against a layerwise variant of AutoFT that learns per-layer learning rates, weight decays, and L1/L2 norms and regularization terms in Tab. 10. The layerwise variant of AutoFT results in worse ID and OOD performance, suggesting that an overly expressive hyperparameter space overfits to the validation set.
Appendix B Experimental Details
B.1 Toy Experiment Details
We provide further details for the toy experiment in Sec. 4.3 and Figure 3. The prior distribution is a unit Gaussian, i.e. it has variance $[{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},\ldots]$. Training data $D_{\text{tr}}$ is drawn from a distribution with variance $[{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},\ldots]$, while the validation data $D_{\text{val}}$ is drawn from a distribution with variance $[{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},{\color[rgb]{0.72,0,0}10^{-3}},{\color[rgb]{0.1,0.0499999999999999,0.96}1.0},\ldots]$. We fit the parameters of a Gaussian distribution $q_{\theta}$ to the training data using a variational Bayes objective by minimizing the following loss:
B.2 Datasets
@@
  • CIFAR-10 [49] contains 60,000 images across 10 classes. We use CIFAR-10 for fine-tuning, 100 examples from CIFAR-10-C for validation, and the CIFAR-10.1 [75, 84] and CIFAR-10.2 [64] as OOD test sets.
@@
  • ImageNet [21] contains over a million images in 1000 categories. We use ImageNet as our ID distribution, 15000 examples from ImageNet-C for validation, and five ImageNet variations for the OOD datasets following prior works [71, 92, 50, 30]: ImageNet-V2 [76], ImageNet-R [37], ImageNet-A [38], ImageNet-Sketch [88], and ObjectNet [7].
@@
  • WILDS-iWildCam [9, 47, 79] is an animal camera trap image classification dataset where the label $y$ is one of 182 animal species. The ID and OOD datasets consist of photos from disjoint sets of camera traps, making the images differ in camera specifications and attributes such as background and lighting. We use the official splits from Koh et al. [47] and use the ID train set for fine-tuning, the OOD validation set for hyperparameter optimization, and the OOD test set for evaluation.
@@
  • WILDS-FMoW [16, 47, 79] contains remote sensing imagery from satellites. Each image is to be classified into one among 62 categories, including labels like “impoverished settlement” and “hospital.” The ID and OOD datasets differ in year of acquisition and geographic location. We use the official splits from Koh et al. [47] and use the ID train set for fine-tuning, the OOD validation set for hyperparameter optimization, and the OOD test set for evaluation.
@@
In all of the transfer learning datasets described below, we use a subset of the ID validation set for hyperparameter optimization. In other words, we do not use an external “OOD” set.
@@
  • Caltech101 [25] contains images of objects from 101 different categories, including “dragonfly,” “grand piano,” and “saxophone.”
@@
  • StanfordCars [48] features a collection of car images categorized by model, make, and year, where the task is to classify them into one of 196 types, such as “Ford Mustang Convertible 1967” or “Toyota Prius Hatchback 2009.”
@@
  • Flowers102 [87] consists of flower images from the UK, with the objective of classifying each image into one of 102 species, such as “oxeye daisy” or “hibiscus.”
@@
  • PatchCamelyon [87] provides digital pathology images for binary classification, with the goal of identifying metastatic tumor tissues.
@@
  • Rendered SST2 [72] is a dataset for optical character recognition, where the task is to classify text sentiment as “positive” or “negative.”
B.3 Training Details
B.3.1 Baselines
We closely follow the training details in Goyal et al. [30]. For all datasets excluding ImageNet, we use a batch size of 256 and conduct a hyperparameter sweep across five learning rates $\{10^{-6},10^{-5},\dots,10^{-2}\}$ and five weight decay values in the range $\{0.0,0.1,\dots,0.4\}$. On ImageNet, we use a larger batch size of 512 and perform a hyperparameter sweep three learning rates $\{10^{-6},10^{-5},10^{-4}\}$ and two weight decays $\{0,0.1\}$. For L2-SP, we tune the weight of the regularization term $\lambda\in\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. We select baseline hyperparameters based on ID validation performance. For datasets without a standard validation set, we split the training data into an 80:20 ratio to create one.
In the $k$ few-shot setting with $k\in\{4,16,32\}$, we select $k$ training and $k$ validation examples from each class. To account for variance from the small training and validation sets, we average over 50 runs.
B.3.2 AutoFT
Hyperparameter search space. As described in Sec. 4, AutoFT learns weights for nine different losses on a log-uniform range $[10^{-4},10]$. AutoFT additionally searches for learning rate in the log-uniform range $[10^{-2}\cdot\eta^{*},10^{2}\cdot\eta^{*}]$, where $\eta^{*}$ is the conventional learning rate used in prior works on fine-tuning [92, 50, 30], and weight decay values in the log-uniform range $[0.0,1.0]$.
Optuna settings. Below, we summarize the number of inner-loop gradient steps, Optuna trials, and validation set size for each dataset. All of these values were selected based on performance on an ID validation set:
Dataset Inner Steps Trials $N_{\textrm{val}}$
iWildCam 10 500 1000
FMoW 10 500 1000
CIFAR-10 10 100 100
Flowers 50 500 500
Cars 50 500 1000
ImageNet 100 500 1000
Few-shot classification. For the $k$ few-shot classification setting, where $k\in\{4,16,32\}$, we use a $k$-shot validation set for hyperparameter optimization. On all $k$-shot SST2 and PatchCamelyon experiments, we run with 10 inner-loop gradient steps and 50 Optuna trials. We run all ImageNet few-shot experiments with 100 Optuna trials and 5, 20, and 50 inner steps for 4-shot, 16-shot, and 32-shot classification, respectively. Due to increased variance from the very small training and validation sets in few-shot binary classification (e.g., SST2 and PatchCamelyon), we select the single best set of hyperparameters across 5 separate AutoFT runs, based on performance on the $k$-shot validation set.
Transfer learning. We note that the transfer learning experiments on Flowers102 and Stanford Cars do not use OOD data for hyperparameter optimization; here, AutoFT optimizes hyperparameters on a held-out ID validation set.
WILDS-iWildCam and WILDS-FMoW state-of-the-art (SoTA). For the iWildCam and FMoW SoTA results in Tab. 1, we fine-tune the ViT-L/14@336px model with loss weights learned on the smaller ViT-B/16 backbone with AutoFT. Applying AutoFT directly to the ViT-L/14@336px model may improve performance further, although at the cost of more compute.

### 07 Mar 2024 v2

@@ B.1.2 Few-Shot Classification
Table 10: AutoFT can lead to improvements in IID transfer learning. In addition to enhancing OOD generalization, AutoFT can improve performance in IID transfer learning settings when using an ID validation set to learn the objective and hyperparameters..
B.2 CIFAR-10 Distribution Shifts
@@
Table 11: Evaluation on CIFAR-10 distribution shifts. AutoFT outperforms fine-tuning by $2.2\%$ on CIFAR-10.2 and by $1.4\%$ on CIFAR-10.1, using only 100 samples from CIFAR-10-C. AutoFT additionally outperforms WiSE-FT with weight ensembling.
AutoFT also outperforms prior robust fine-tuning approaches on the subtle CIFAR-10.1 and CIFAR-10.2 distribution shifts. Here, AutoFT uses the CLIP ViT-L/14 backbone, following Wortsman et al. (2022b), and $100$ examples from CIFAR-10-C to learn the fine-tuning objective and hyperparameters.
Appendix C Additional Experiments
C.1 Expressivity of Hyperparameter Space
@@
Table 12: Expressivity of the AutoFT search space. AutoFT results in better performance than its layerwise variant with per-layer learning rates, weight decays, and L1/L2 distances to pretrained parameters. Evaluation is on the iWildCam benchmark. This suggests that more expressive hyperparameter spaces overfit to the validation set, degrading performance.
@@
We investigate whether more expressive hyperparameter search spaces result in better performance. We compare AutoFT against a layerwise variant of AutoFT that learns per-layer learning rates, weight decays, and L1/L2 norms and regularization terms in Table 12. The layerwise variant of AutoFT results in worse ID and OOD performance, suggesting that an overly expressive search space leads to overfitting of the small validation set.
Appendix D Experimental Details
D.1 Training Protocol
We closely follow the training details of Goyal et al. (2022a) and Wortsman et al. (2022b). All methods fine-tune models with an AdamW optimizer, cosine learning rate scheduler, and a batch size of 512 for ImageNet and 256 for all other datasets. All baseline hyperparameters, such as learning rate, weight decay, and warmup length, are tuned through grid search. All methods, including AutoFT, perform early stopping based on in-distribution (ID) validation accuracy. We provide a comprehensive breakdown of the hyperparameter sweeps in the supplementary material. We emphasize that none of these methods, including AutoFT, observe any of the test OOD distributions during training. Finally, we report metrics averaged over 5 runs with 95% confidence intervals.
D.2 Baselines
We follow the training details in Goyal et al. (2022a). For all datasets excluding ImageNet, we use a batch size of 256 and conduct a hyperparameter sweep across five learning rates $\{10^{-6},10^{-5},\dots,10^{-2}\}$ and five weight decay values in the range $\{0.0,0.1,\dots,0.4\}$. On ImageNet, we use a larger batch size of 512 and perform a hyperparameter sweep over three learning rates $\{10^{-6},10^{-5},10^{-4}\}$ and two weight decays $\{0,0.1\}$. For L2-SP, we tune the weight of the regularization term $\lambda\in\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. We select baseline hyperparameters based on ID validation performance. For datasets without a standard validation set, we split the training data into an 80:20 ratio to create one.
D.3 AutoFT
Hyperparameter search space. As described in Section 4, AutoFT learns weights for nine different losses on a log-uniform range $[10^{-4},10]$. AutoFT additionally searches for learning rate in the log-uniform range $[10^{-2}\cdot\eta^{*},10^{2}\cdot\eta^{*}]$, where $\eta^{*}$ is the conventional learning rate used in prior works on fine-tuning (Wortsman et al., 2022b; Kumar et al., 2022a; Goyal et al., 2022a), weight decay values in the log-uniform range $[0.0,1.0]$, and random seeds in the range $[0,100]$. The layerwise variant of AutoFT learns per-layer learning rates and weight decay regularization terms, each sampled from the same respective range.
WILDS-iWildCam and WILDS-FMoW state-of-the-art (SoTA). For the iWildCam and FMoW SoTA results in Table 2, we fine-tune the ViT-L/14@336px model with loss weights learned on the smaller ViT-B/16 backbone with AutoFT. Applying AutoFT directly to the ViT-L/14@336px model may improve performance further, although at the cost of more compute.
Validation sets. For the iWildCam and FMoW experiments, we use $1000$ examples from the official WILDS OOD validation splits from Koh et al. (2021). For ImageNet, we use $1000$ examples from ImageNet-C. For CIFAR-10, we use $100$ examples from CIFAR-10-C. For the few-shot experiments, we use a randomly sampled $k$-shot ID validation set, and average over 50 runs to account for variance in the fine-tuning and validation sets. For the transfer learning experiments, we do not use OOD data. We randomly partition the ID validation sets into two validation sets: one for hyperparameter optimization, and the other for early stopping during final fine-tuning. We use $100,200,400$ examples for Caltech-101, Stanford Cars, and Flowers-102, respectively.
Hyper-hyperparameters. We list hyper-hyperparameters used in Table 1. Few-shot hyper-hyperparameters are listed in Table 13. We run AutoFT with $500$ outer-loop trials on all non-few shot experiments, which is the largest number of trials that can be run with a reasonable amount of compute. We use grid-search to tune the number of inner steps and validation set size and select hyper-hyperparameters based on ID validation performance of the final fine-tuned model. On iWildCam, FMoW, and CIFAR-10, we tune over $\{10,50,100\}$ inner steps and $\{100,1000,5000\}$ validation examples. On ImageNet, we tune over $\{10,50\}$ inner steps and $\{1000,5000\}$ validation examples. On Caltech-101, Stanford Cars, and Flowers-102, we tune over $\{10,50\}$ inner steps. For the few-shot binary classification experiments, we run AutoFT with $50$ outer-loop evaluations and tune over $\{5,10,20\}$ inner steps. For the few-shot ImageNet experiments, we run AutoFT with $100$ outer-loop evaluations and tune over $\{5,20,50\}$ inner steps.
Dataset Steps Trials
SST-2 10 50
PatchCamelyon 10 50
ImageNet-4 5 100
ImageNet-16 20 100
ImageNet-32 50 100
Table 13: AutoFT training settings for few-shot classification experiments.
Our empirical results show that a small number of inner loop steps is effective for identifying suitable hyperparameters. This is also a necessity considering computational constraints. Our method strikes a balance between practical feasibility and reliable hyperparameter selection.
Few-shot classification. In the $k$-shot setting with $k\in\{4,16,32\}$, we select $k$ training and $k$ validation examples from each class. To account for variance from the small training and validation sets, for each run of AutoFT, we repeat hyperparameter optimization $5$ times and select the best set of hyperparameters based on ID validation performance. We then average results over $50$ runs.
Transferability of learned hyperparameters experiment details. The OOD distributions for fine-tuning on CIFAR-10 are CIFAR-10.1 and CIFAR-10.2. The OOD distributions for fine-tuning on MNIST are MNIST-C, rotated MNIST, colored MNIST, and EMNIST.
D.4 Datasets
@@
  • CIFAR-10 (Krizhevsky et al., 2009) contains 60,000 images across 10 classes. We use CIFAR-10 for fine-tuning, 100 examples from CIFAR-10-C for validation, and the CIFAR-10.1 (Recht et al., 2018; Torralba et al., 2008) and CIFAR-10.2 (Lu et al., 2020) as OOD test sets.
@@
  • ImageNet (Deng et al., 2009) contains over a million images in 1000 categories. We use ImageNet as our ID distribution, 15000 examples from ImageNet-C for validation, and five ImageNet variations for the OOD datasets following prior works (Radford et al., 2021a; Wortsman et al., 2022b; Kumar et al., 2022a; Goyal et al., 2022a): ImageNet-V2 (Recht et al., 2019), ImageNet-R (Hendrycks et al., 2021a), ImageNet-A (Hendrycks et al., 2021b), ImageNet-Sketch (Wang et al., 2019), and ObjectNet (Barbu et al., 2019).
@@
  • WILDS-iWildCam (Beery et al., 2021; Koh et al., 2021; Sagawa et al., 2022) is an animal camera trap image classification dataset where the label $y$ is one of 182 animal species. The ID and OOD datasets consist of photos from disjoint sets of camera traps, making the images differ in camera specifications and attributes such as background and lighting. We use the official splits from Koh et al. (2021) and use the ID train set for fine-tuning, the OOD validation set for hyperparameter optimization, and the OOD test set for evaluation.
@@
  • WILDS-FMoW (Christie et al., 2018; Koh et al., 2021; Sagawa et al., 2022) contains remote sensing imagery from satellites. Each image is to be classified into one among 62 categories, including labels like “impoverished settlement” and “hospital.” The ID and OOD datasets differ in year of acquisition and geographic location. We use the official splits from Koh et al. (2021) and use the ID train set for fine-tuning, the OOD validation set for hyperparameter optimization, and the OOD test set for evaluation.
@@
In all of the transfer learning datasets described below, we use a subset of the ID validation set for hyperparameter optimization. In other words, we do not use an “OOD” validation set.
@@
  • Caltech101 (Fei-Fei et al., 2004) contains images of objects from 101 different categories, including “dragonfly,” “grand piano,” and “saxophone.”
@@
  • StanfordCars (Krause et al., 2013) features a collection of car images categorized by model, make, and year, where the task is to classify them into one of 196 types, such as “Ford Mustang Convertible 1967” or “Toyota Prius Hatchback 2009.”
@@
  • Flowers102 (Veeling et al., 2018) consists of flower images from the UK, with the objective of classifying each image into one of 102 species, such as “oxeye daisy” or “hibiscus.”
@@
  • PatchCamelyon (Veeling et al., 2018) provides digital pathology images for binary classification, with the goal of identifying metastatic tumor tissues.
@@
  • Rendered SST2 (Radford et al., 2021b) is a dataset for optical character recognition, where the task is to classify text sentiment as “positive” or “negative.”